A suture is stretched to 10%. When the stress is released it recovered 35 % of its strain after 5 hr. What is the retardation time in hours. What is the amount of strain recovered (in %) after 6 hrs of its original value.
1
Expert's answer
2020-03-10T11:36:58-0400
"\frac{l}{l_0}(t)=1-e^{-\frac{t}{\tau}}"

1)


"\frac{l}{l_0}(5)=1-e^{-\frac{5}{\tau}}=0.35"

"\tau=11.6\ h"

2)


"\frac{l}{l_0}(6)=1-e^{-\frac{6}{11.6}}=0.40"

The amount of strain recovered after 6 hrs is 40% of its original value.